The width of a rectangular yard is x feet. If 300 feet of fencing is needed to enclse the yard, which of the following represents the length of the yard, in feet, and why?
A. 300-x
B. 300-2x
C. 150-x
D. 150-2x

Answer by flame8855(424)

since the perimeter is = 2 ( l+w)
so 2 ( L + W ) = 300
L+w = 150
L= 150-w
L= 150 -x
so the answer is C